If it takes 6 workers one hour to put together one car, how long would it take 9 workers to produce the same
car?

Answer by stanbon(75887)

If it takes 6 workers one hour to put together one car, how long would it take 9 workers to produce the same
car?
====
# of workers and hours are indirectly related:
w = k/h
---
Solve for "k" using "it takes 6 workers one hour to put together one car".
6 = k/1
k = 6
----
Equation:
w = 6/h
---
How long would it take 9 workers to produce the same car?
9 = 6/h
h = 6/9
h = 2/3 hr = 40 minutes
==============================
Cheers,
Stan H.
==============